OpenStudy (smurfy14):

[(x)/(x+2)]-[(2)/(2x-1)]=(1/5) please solve thanks!

OpenStudy (immanuelv):

one of the solution is -3/4 other one is 3

OpenStudy (smurfy14):

how though?

OpenStudy (anonymous):

the solution is in the attachment

OpenStudy (smurfy14):

thanks!! do you know how to get -3/4 like that iman guy was talking about?

OpenStudy (anonymous):

there is a mistake. Im trying to fix it

OpenStudy (immanuelv):

do you have a graphing calculator?

OpenStudy (smurfy14):

no but im not allowed to use it on my hw anyways

OpenStudy (immanuelv):

ok so what u need to do is plug in 1/5 to the y then solve for x!

OpenStudy (smurfy14):

emunrradtvamg are you talking about the -1 being positive? if so i got it. and imman there is no y?

OpenStudy (immanuelv):

well 1/5 is basically the y sorry!

OpenStudy (anonymous):

This is where i got

OpenStudy (anonymous):

and finally

OpenStudy (anonymous):

im i right?

OpenStudy (smurfy14):

i dont think thats right lol but the very first one you attached is right. now i just need to figure out how to get -3/4 as the second answer. The correct answer is 3 and -3/4 i just dont know how to show the work for the second answer.

OpenStudy (anonymous):

actually the very first attachment is wrong

OpenStudy (smurfy14):

well you messed up a bit at the end but in the end it is still 3

OpenStudy (smurfy14):

5x-10=2x-1 3x=9 x=3

OpenStudy (anonymous):

Ok. I tried lol

OpenStudy (anonymous):

too long to be solved through openstudy

OpenStudy (smurfy14):

thanks very much! do you how to get -3/4 for the second answer?

OpenStudy (radar):

I too, am having trouble following the work on the second attachment. Although 3 is a correct solution.

OpenStudy (smurfy14):

do you know how to get -3/4 radar?

OpenStudy (radar):

Well the quadratic may have two roots, the 3 and another. I am hung up on the second line of the attachment, where he put the numerator over the LCD and got x(2x+2) I keep getting x(2x-1) for the first fraction. I will look into it further and post here later. Right now I got to put my dogs up.

OpenStudy (radar):

But first lets see if -3/4 is an answer -3/4 2 1 ------ - --------- = ---- -3/4 +2 2(-3/4)-1 5 -3/4 2 1 ------ - ---------- = ---- 5/4 -5/2 5 -3/5 - (-4/5 = 1/5 It checks as an answer

OpenStudy (radar):

The solution to that quadratic does not come out to 3 or -3/4. the results of\[x=1/16(3\pm 3\sqrt{65}\] is an irrational number not the integer 3, or the fraction -3/4 something is wrong (maybe with me lol)